Hi, Im a little confused with this problem. I tried to figure it out, but do not know what would be the right answer.
Solve by the elimination method.
5x+6y=5
10x+12y=10
what is the solution of the system?
( I came up with - The solution is (10,10).
but could the answer be that there are infinetely many solutions, or that there is no solution. can you please help. thanks

Answer by Theo(13342)

5x+6y=5
10x+12y=10
-----
multiply first equation by 2 to get:
10x + 12y = 10
10x + 12y = 10
-----
subtract second equation from first to get:
0 + 0 = 0
-----
I believe the answer is that this is an identity and that any value of x and y that satisfied one equation would also satisfy the other.
-----
if you graph both these equations, you would probably see that they are the same line.
-----
first equation:
5x + 6y = 5
solving for y we get:
6y -5x + 5
y = -(5/6)x + (5/6)
-----
second equation:
10x + 12y = 10
solving for y we get:
12y = -10x + 10
y = (-10/12)x + (10/12)
which can be reduced to:
y = (-5/6)x + (5/6)
-----
the equations are identical after simplification.